# Clarifying the effect of mean subtraction on Dynamic Mode Decomposition

### v5

@@ A.1 The Koopman mode factorization of $\mathbf{Z}$
Specifically, if we define $\mathbf{C}$ to be the column scaling of $\tilde{\mathbf{C}}$ (3.8) by the eigen-coordinates of the initial condition $\boldsymbol{\mathbf{\upsilon}}_{1}$,
@@
Proof A.2.
We begin with the definition of $\mathbf{Z}$ in Eq. 4:
$\boldsymbol{\mathbf{z}}_{j}~{}=~{}\boldsymbol{\mathbf{\psi}}(\Gamma^{j}(\boldsymbol{\mathbf{\upsilon}}_{1})).$
Using Definition 3.8, if we write $\boldsymbol{\mathbf{\psi}}$ in terms of the eigenfunctions $\boldsymbol{\mathbf{\phi}}$, then, the potentially nonlinear map $\Gamma^{j}$ can be replaced with the Koopman operator (Definitions 3.1 and 14).
$\underbrace{\boldsymbol{\mathbf{\psi}}}(\Gamma^{j}(\boldsymbol{\mathbf{\upsilon}}_{1}))~{}=~{}\tilde{\mathbf{C}}\left(\underbrace{\boldsymbol{\mathbf{\phi}}\circ\Gamma^{j}}\right)(\boldsymbol{\mathbf{\upsilon}}_{1})~{}=~{}\tilde{\mathbf{C}}\left(U^{j}\circ\boldsymbol{\mathbf{\phi}}\right)(\boldsymbol{\mathbf{\upsilon}}_{1}).$
If we unpack each component of $\left(U^{j}\circ\boldsymbol{\mathbf{\phi}}\right)(\boldsymbol{\mathbf{\upsilon}}_{1})$ and invoke Definition 3.2,
$\left(U^{j}\circ\boldsymbol{\mathbf{\phi}}\right)(\boldsymbol{\mathbf{\upsilon}}_{1})~{}=~{}\left[\left(\underbrace{U^{j}\circ\phi_{i}}\right)(\boldsymbol{\mathbf{\upsilon}}_{1})\right]_{i=1}^{r}~{}=~{}\left[\lambda_{i}^{j}\phi_{i}(\boldsymbol{\mathbf{\upsilon}}_{1})\right]_{i=1}^{r},$
@@
$\tilde{\mathbf{C}}\left[\lambda_{i}^{j}\underbrace{\phi_{i}(\boldsymbol{\mathbf{\upsilon}}_{1})}\right]_{i=1}^{r}~{}=~{}\underbrace{\tilde{\mathbf{C}}~{}\operatorname{diag}[\boldsymbol{\mathbf{\phi}}(\boldsymbol{\mathbf{\upsilon}}_{1})]}\left[\lambda_{i}^{j}\right]_{i=1}^{r}~{}=~{}\mathbf{C}\left[\lambda_{i}^{j}\right]_{i=1}^{r}.$
Compiling this expression for each column of $\mathbf{Z}$ gives Eq. 24. Lastly, Definition 3.10 and Eq. 23, when taken together, readily yield Eq. 25.
Remark A.3.
@@
A.2 An analytically useful re-phrasal of Koopman invariance
Proposition A.4.
Suppose $\boldsymbol{\mathbf{\psi}}$ lies in the non-redundant span of distinct KEFs. Then, Koopman invariance of $\boldsymbol{\mathbf{\psi}}$ is equivalent to the columns of $\tilde{\mathbf{C}}$ being linearly independent.
@@ A.2 An analytically useful re-phrasal of Koopman invariance
Under the additional condition of a spectrally informative $\boldsymbol{\mathbf{\upsilon}}_{1}$, Koopman invariance can be inferred, alternatively, from the matrix $\mathbf{C}$:
@@
Proof A.5.
We will focus on establishing Eq. 29 since it readily leads [^16] When combined with Definition 3.10 and Eq. 23. to Eq. 30.
We begin with the backward implication. Consider the $m\times m$ matrix,
@@
where $\mathbf{\Lambda}$ is from Eq. 27. We will show that $U\circ\boldsymbol{\mathbf{\psi}}~{}=~{}\mathbf{A}\boldsymbol{\mathbf{\psi}}$, by expanding $\boldsymbol{\mathbf{\psi}}$ on the right-hand side using Definition 3.8 and, then, unpacking $\mathbf{A}$:
$\mathbf{A}\underbrace{\boldsymbol{\mathbf{\psi}}}~{}=~{}\underbrace{\mathbf{A}}\tilde{\mathbf{C}}\boldsymbol{\mathbf{\phi}}~{}=~{}(\tilde{\mathbf{C}}\mathbf{\Lambda}\tilde{\mathbf{C}}^{\dagger})\tilde{\mathbf{C}}\boldsymbol{\mathbf{\phi}}.$
Since $\tilde{\mathbf{C}}$ has full column rank, $\tilde{\mathbf{C}}^{\dagger}\tilde{\mathbf{C}}=\mathbf{I}$. Using this identity lets us apply Eqs. 27 and 14 simultaneously to bring in the Koopman operator $U$:
$\tilde{\mathbf{C}}\mathbf{\Lambda}\underbrace{\tilde{\mathbf{C}}^{\dagger}\tilde{\mathbf{C}}}\boldsymbol{\mathbf{\phi}}~{}=~{}\tilde{\mathbf{C}}\underbrace{\mathbf{\Lambda}\boldsymbol{\mathbf{\phi}}}~{}=~{}\tilde{\mathbf{C}}~{}\left(U\circ\boldsymbol{\mathbf{\phi}}\right).$
Finally, we can use the linearity of $U$ to link $\tilde{\mathbf{C}}$ and $\boldsymbol{\mathbf{\phi}}$:
$\underbrace{\tilde{\mathbf{C}}~{}U}\circ\boldsymbol{\mathbf{\phi}}~{}=~{}U\circ\underbrace{\tilde{\mathbf{C}}\boldsymbol{\mathbf{\phi}}}~{}=~{}U\circ\boldsymbol{\mathbf{\psi}}.$
For the forward implication, recall that a Koopman invariant $\boldsymbol{\mathbf{\psi}}$ implies, by Definition 4.7, the existence of a matrix $\mathbf{A}$ such that
@@
Now, for the left-hand side, we can simply reverse the steps taken towards the conclusion of the forward implication- Invoke Definition 3.8, follow it up with the linearity of $U$ and conclude by applying Eqs. 27 and 14 simultaneously:
$U\circ\underbrace{\boldsymbol{\mathbf{\psi}}}~{}=~{}\underbrace{U\circ\tilde{\mathbf{C}}}\boldsymbol{\mathbf{\phi}}~{}=~{}\tilde{\mathbf{C}}\underbrace{U\circ\boldsymbol{\mathbf{\phi}}}~{}=~{}\tilde{\mathbf{C}}\mathbf{\Lambda}\boldsymbol{\mathbf{\phi}}.$
For the right-hand side, simply apply Definition 3.8.
$\mathbf{A}\underbrace{\boldsymbol{\mathbf{\psi}}}~{}=~{}\mathbf{A}\tilde{\mathbf{C}}\boldsymbol{\mathbf{\phi}}.$
@@
Appendix B The size of a Vandermonde matrix can determine its rank
@@
Lemma B.1 (Node distinctness lets size dictate rank).
@@
Appendix C On approximating the Koopman eigen-values via DMD
Under the assumptions (3.8) and (3.10), we prove that linearly consistent and well-sampled training data produces DMD eigen-values, $\{\,\hat{\lambda}_{i}\,\}_{i=1}^{n}$, that contain the true Koopman eigen-values, $\{\,\lambda_{i}\,\}_{i=1}^{r}$. Furthermore, the latter are characterized by the corresponding DMD modes being non-zero.
@@ Appendix C On approximating the Koopman eigen-values via DMD
Proposition C.1.
Suppose (3.8) is true. If the DMD problem is well-sampled, then, the last snapshot, $\boldsymbol{\mathbf{z}}_{n+1}$, is a linear combination of the preceding snapshots.
@@
Proof C.2.
We can simply recast the arguments used in Theorem 1 of [18], using the machinery developed so far. According to Eqs. 33 and 32, the condition $n\geq r$ ensures that the last column of $\mathbf{\Theta}$ lies in the span of $\mathbf{\Theta}_{1}$ i.e.,
@@
If we pre-multiply both sides with $\mathbf{C}$, then, we can apply Eqs. 24 and 28 to reach the desired conclusion.
@@
Proposition C.3.
Suppose (3.8) and (3.10) are true. If $(\mathbf{X},\mathbf{Y})$ is linearly consistent and the DMD problem is well-sampled, then, the columns of $\tilde{\mathbf{C}}$ are linearly independent.
@@
Proof C.4.
Linear consistency of $(\mathbf{X},\mathbf{Y})$, by Definition 3.11, implies there exists a matrix $\mathbf{A}$ such that $\mathbf{A}\mathbf{X}=\mathbf{Y}$. We may expand the data matrices using Eq. 28 to obtain:
$\mathbf{A}\underbrace{\mathbf{X}}=\underbrace{\mathbf{Y}}\iff(\mathbf{A}\mathbf{C})\mathbf{\Theta}_{1}=(\mathbf{C}\mathbf{\Lambda})\mathbf{\Theta}_{1}.$
@@
$\mathbf{A}\underbrace{\mathbf{C}}\mathbf{\Theta}_{1}{}$ $=~{}\mathbf{A}\tilde{\mathbf{C}}\operatorname{diag}[\boldsymbol{\mathbf{\phi}}(\boldsymbol{\mathbf{\upsilon}}_{1})]\mathbf{\Theta}_{1},$
$\underbrace{\mathbf{C}}\mathbf{\Lambda}\mathbf{\Theta}_{1}~{}=~{}\tilde{\mathbf{C}}~{}\underbrace{\operatorname{diag}[\boldsymbol{\mathbf{\phi}}(\boldsymbol{\mathbf{\upsilon}}_{1})]\mathbf{\Lambda}}\mathbf{\Theta}_{1}{}$ $=~{}(\tilde{\mathbf{C}}\mathbf{\Lambda})~{}\operatorname{diag}[\boldsymbol{\mathbf{\phi}}(\boldsymbol{\mathbf{\upsilon}}_{1})]\mathbf{\Theta}_{1}.$
Combining these results, we have:
The matrix $\operatorname{diag}[\boldsymbol{\mathbf{\phi}}(\boldsymbol{\mathbf{\upsilon}}_{1})]\mathbf{\Theta}_{1}$ has full row rank. To see this, we may first use $\boldsymbol{\mathbf{\upsilon}}_{1}$ being spectrally informative (3.10) alongside Eq. 23 to infer that $\operatorname{diag}[\boldsymbol{\mathbf{\phi}}(\boldsymbol{\mathbf{\upsilon}}_{1})]$ is invertible. Then, we can consider the $r\times n$ Vandermonde matrix $\mathbf{\Theta}_{1}$. Since it has distinct nodes, well-sampling $(n\geq r)$ lets us use Eq. 32 to conclude that $\mathbf{\Theta}_{1}$ has linearly independent rows. Since the product of two matrices with full row rank will also have full row rank, we find that the rows of $\operatorname{diag}[\boldsymbol{\mathbf{\phi}}(\boldsymbol{\mathbf{\upsilon}}_{1})]\mathbf{\Theta}_{1}$ are linearly independent. Hence, Eq. 34 is equivalent to the following identity:
@@
By Eq. 27, $\mathbf{\Lambda}$ is a diagonal matrix constructed by collecting the Koopman eigen-values, $\{\,\lambda_{i}\,\}_{i=1}^{r}$, which, according to (3.8), are $r$ distinct complex numbers. The resulting re-interpretation of Eq. 35- the columns of $\tilde{\mathbf{C}}$ are eigen-vectors of $\mathbf{A}$ corresponding to distinct eigen-values- yields the desired conclusion.
@@
Proof C.5 (Proof of Theorem 3.4).
We preface the proof by establishing that $\mathbf{C}$ has full column rank. Firstly, Proposition C.3 tells us that the columns of $\tilde{\mathbf{C}}$ are linearly independent. Building on this, we only need apply (23) and (3.10) to show that $\mathbf{C}$ also has full column rank.
Now, consider the least-squares interpretation of Eq. 6:
@@
Since $n\geq r$, Proposition C.1 tells us that the first branch is active. Hence,
@@
Now, we can expand both $\mathbf{X}$ and $\mathbf{Y}$ according to Eq. 28, and then safely discard $\mathbf{C}$ owing to its full column rank:
$\underbrace{\mathbf{X}}\mathbf{T}^{*}-\underbrace{\mathbf{Y}}~{}=~{}\mathbf{0}\iff\underbrace{\mathbf{C}}\left(\mathbf{\Theta}_{1}\mathbf{T}^{*}-\mathbf{\Lambda}\mathbf{\Theta}_{1}\right)~{}=~{}\mathbf{0}\iff\mathbf{\Theta}_{1}\mathbf{T}^{*}-\mathbf{\Lambda}\mathbf{\Theta}_{1}~{}=~{}\mathbf{0}.$
Since $\mathbf{\Lambda}$ is a diagonal matrix, the final identity,

### v6

@@ B.1 The Koopman mode factorization of $\mathbf{Z}$
Specifically, if we define $\mathbf{C}$ to be the column scaling of $\tilde{\mathbf{C}}$ by the eigen-coordinates of the initial condition $\boldsymbol{\mathbf{\upsilon}}_{1}$,
@@
Proof
Deducing the Koopman mode factorization of $\mathbf{Z}$
Recall the construction of $\mathbf{Z}$ from Eq. 4:
$\mathbf{Z}~{}=~{}\left[\boldsymbol{\mathbf{z}}_{j}\right]_{j=0}^{n}~{}=~{}\left[\boldsymbol{\mathbf{\psi}}\left(\Gamma^{j}\left(\boldsymbol{\mathbf{\upsilon}}_{1}\right)\right)\right]_{j=0}^{n}.$
Here, for the sake of convenience, we have numbered the columns of $\mathbf{Z}$ from 00 to $n$, instead of the original choice of $1$ to $n+1$. We can begin by using Eq. 20 to write $\boldsymbol{\mathbf{\psi}}$ in terms of the eigenfunctions $\boldsymbol{\mathbf{\phi}}$ and, then, leverage Eq. 23 to replace the potentially nonlinear map $\Gamma^{j}$ with the Koopman operator.
${\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\boldsymbol{\mathbf{\psi}}}(\Gamma^{j}(\boldsymbol{\mathbf{\upsilon}}_{1}))~{}=~{}\tilde{\mathbf{C}}\left({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\boldsymbol{\mathbf{\phi}}\circ\Gamma^{j}}\right)(\boldsymbol{\mathbf{\upsilon}}_{1})~{}=~{}\tilde{\mathbf{C}}\left(U^{j}\circ\boldsymbol{\mathbf{\phi}}\right)(\boldsymbol{\mathbf{\upsilon}}_{1}).$
Now, if we unpack each component of $\left(U^{j}\circ\boldsymbol{\mathbf{\phi}}\right)(\boldsymbol{\mathbf{\upsilon}}_{1})$ and invoke the defining property of a Koopman eigen-function (3.2),
$\left(U^{j}\circ\boldsymbol{\mathbf{\phi}}\right)(\boldsymbol{\mathbf{\upsilon}}_{1})~{}=~{}\left[\left({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}U^{j}\circ\phi_{i}}\right)(\boldsymbol{\mathbf{\upsilon}}_{1})\right]_{i=1}^{r}~{}=~{}\left[\lambda_{i}^{j}\phi_{i}(\boldsymbol{\mathbf{\upsilon}}_{1})\right]_{i=1}^{r},$
@@
$\tilde{\mathbf{C}}\left[\lambda_{i}^{j}{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\phi_{i}(\boldsymbol{\mathbf{\upsilon}}_{1})}\right]_{i=1}^{r}~{}=~{}{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\tilde{\mathbf{C}}~{}\operatorname{diag}[\boldsymbol{\mathbf{\phi}}(\boldsymbol{\mathbf{\upsilon}}_{1})]}\left[\lambda_{i}^{j}\right]_{i=1}^{r}~{}=~{}\mathbf{C}\left[\lambda_{i}^{j}\right]_{i=1}^{r}.$
Compiling this expression for each column of $\mathbf{Z}$ gives Eq. 36.
Connecting spectral informativeness to the absence of redundancies
According to Eq. 21, every column of $\tilde{\mathbf{C}}$ is non-zero.
When $\boldsymbol{\mathbf{\upsilon}}_{1}$ is spectrally informative, each of the $r$ KEFs, $\{\,\phi_{i}\,\}_{i=1}^{r}$, evaluates to a non-zero value at $\boldsymbol{\mathbf{\upsilon}}_{1}$. Since these same non-zero values scale the columns of $\tilde{\mathbf{C}}$, in Eq. 35, to produce $\mathbf{C}$, every column of $\mathbf{C}$ is also non-zero.
On the other hand, if $\boldsymbol{\mathbf{\upsilon}}_{1}$ is not spectrally informative, then, there is a KEF that evaluates to zero at $\boldsymbol{\mathbf{\upsilon}}_{1}$. Hence, we can apply Eq. 35 to deduce that one of the columns in $\mathbf{C}$ is zero.
[…]
Remark B.2.
@@
B.2 An analytically useful re-phrasal of Koopman invariance
Proposition B.3.
Suppose $\boldsymbol{\mathbf{\psi}}$ lies in the non-redundant span of $r$ distinct KEFs. Then, Koopman invariance of $\boldsymbol{\mathbf{\psi}}$ is equivalent to the columns of $\tilde{\mathbf{C}}$ being linearly independent.
@@ B.2 An analytically useful re-phrasal of Koopman invariance
[…]
@@
Proof
We will focus on establishing Eq. 41 because it is readily transmuted by a spectrally informative $\boldsymbol{\mathbf{\upsilon}}_{1}$ into Eq. 42 .
Koopman invariance via $\tilde{\mathbf{C}}$
The backward implication can be proven using the $m\times m$ matrix,
@@
as the guess for the action of the Koopman operator on $\boldsymbol{\mathbf{\psi}}$. In particular, we will show that $U\circ\boldsymbol{\mathbf{\psi}}~{}=~{}\mathbf{A}\boldsymbol{\mathbf{\psi}}$. To begin with, we can expand $\boldsymbol{\mathbf{\psi}}$ on the right-hand side using Eq. 20 and, then, unpack $\mathbf{A}$:
$\mathbf{A}{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\boldsymbol{\mathbf{\psi}}}~{}=~{}{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\mathbf{A}}\tilde{\mathbf{C}}\boldsymbol{\mathbf{\phi}}~{}=~{}(\tilde{\mathbf{C}}\mathbf{\Lambda}\tilde{\mathbf{C}}^{\dagger})\tilde{\mathbf{C}}\boldsymbol{\mathbf{\phi}}.$
Since $\tilde{\mathbf{C}}$ has full column rank, $\tilde{\mathbf{C}}^{\dagger}\tilde{\mathbf{C}}=\mathbf{I}$. Using this identity lets us apply Eqs. 39 and 23 simultaneously to bring in the Koopman operator $U$:
$\tilde{\mathbf{C}}\mathbf{\Lambda}{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\tilde{\mathbf{C}}^{\dagger}\tilde{\mathbf{C}}}\boldsymbol{\mathbf{\phi}}~{}=~{}\tilde{\mathbf{C}}{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\mathbf{\Lambda}\boldsymbol{\mathbf{\phi}}}~{}=~{}\tilde{\mathbf{C}}~{}\left(U\circ\boldsymbol{\mathbf{\phi}}\right).$
Finally, we can use the linearity of $U$ to link $\tilde{\mathbf{C}}$ and $\boldsymbol{\mathbf{\phi}}$, thereby permitting another application of Eq. 20:
${\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\tilde{\mathbf{C}}~{}U}\circ\boldsymbol{\mathbf{\phi}}~{}=~{}U\circ{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\tilde{\mathbf{C}}\boldsymbol{\mathbf{\phi}}}~{}=~{}U\circ\boldsymbol{\mathbf{\psi}}.$
The forward implication can be tackled by starting with the definition of a Koopman invariant dictionary and reversing the arguments used above. Recall that a Koopman invariant $\boldsymbol{\mathbf{\psi}}$ implies, by Definition 4.7, the existence of a matrix $\mathbf{A}$ such that
@@
Now, for the left-hand side, we can simply reverse the steps taken towards the conclusion of the forward implication- Invoke Eq. 20, follow it up with the linearity of $U$ and conclude by applying Eqs. 39 and 23 simultaneously:
$U\circ{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\boldsymbol{\mathbf{\psi}}}~{}=~{}{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}U\circ\tilde{\mathbf{C}}}\boldsymbol{\mathbf{\phi}}~{}=~{}\tilde{\mathbf{C}}{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}U\circ\boldsymbol{\mathbf{\phi}}}~{}=~{}\tilde{\mathbf{C}}\mathbf{\Lambda}\boldsymbol{\mathbf{\phi}}.$
For the right-hand side, simply apply Eq. 20.
$\mathbf{A}{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\boldsymbol{\mathbf{\psi}}}~{}=~{}\mathbf{A}\tilde{\mathbf{C}}\boldsymbol{\mathbf{\phi}}.$
@@
Koopman invariance via $\mathbf{C}$
When the initial condition $\boldsymbol{\mathbf{\upsilon}}_{1}$ is spectrally informative, then, $\mathbf{C}$ has full column rank if and only if $\tilde{\mathbf{C}}$ also possesses full column rank. According to Definition 3.13, $\boldsymbol{\mathbf{\upsilon}}_{1}$ is spectrally informative when each of the $r$ Koopman eigenfunctions, $\{\,\phi_{i}\,\}_{i=1}^{r}$, evaluates to a non-zero value at $\boldsymbol{\mathbf{\upsilon}}_{1}$. Now, Eq. 35 tells us that these same values are diagonally stacked to form $\operatorname{diag}[\boldsymbol{\mathbf{\phi}}(\boldsymbol{\mathbf{\upsilon}}_{1})]$. In other words,
$\boldsymbol{\mathbf{\upsilon}}_{1}~{}\textrm{is spectrally informative}~{}\iff~{}\operatorname{diag}[\boldsymbol{\mathbf{\phi}}(\boldsymbol{\mathbf{\upsilon}}_{1})]~{}\textrm{is invertible.}$.
Furthermore, this matrix is used to construct $\mathbf{C}$ from $\tilde{\mathbf{C}}$:
$\mathbf{C}~{}=~{}\tilde{\mathbf{C}}~{}\operatorname{diag}[\boldsymbol{\mathbf{\phi}}(\boldsymbol{\mathbf{\upsilon}}_{1})].$
Therefore, when $\boldsymbol{\mathbf{\upsilon}}_{1}$ is spectrally informative, we have,
$\tilde{\mathbf{C}}~{}\textrm{has full column rank}~{}\iff~{}\mathbf{C}~{}\textrm{has full column rank.}$.
Taken together with Eq. 41, we see that Eq. 42 holds as well.
[…]
Appendix C The size of a Vandermonde matrix can determine its rank
@@
Lemma C.2 (Node distinctness lets size dictate rank).
@@
Appendix D On approximating the Koopman eigen-values via DMD
Assuming that $\boldsymbol{\mathbf{\psi}}$ lies in the non-redundant span of $r$ distinct KEFs  and the initial condition $\boldsymbol{\mathbf{\upsilon}}_{1}$ is spectrally informative, we prove that linearly consistent and well-sampled training data produces DMD eigen-values, $\{\,\hat{\lambda}_{i}\,\}_{i=1}^{n}$, that contain the true Koopman eigen-values, $\{\,\lambda_{i}\,\}_{i=1}^{r}$. Furthermore, the latter are characterized by the corresponding DMD modes being non-zero.
@@ Appendix D On approximating the Koopman eigen-values via DMD
Proposition D.1.
Suppose $\boldsymbol{\mathbf{\psi}}$ lies in the non-redundant span of $r$ distinct KEFs. If the DMD problem is well-sampled, then, the last snapshot, $\boldsymbol{\mathbf{z}}_{n+1}$, is a linear combination of the preceding snapshots.
@@
Proof D.2.
We just need to recast the arguments used in Theorem 1 of [19], using the machinery developed so far. According to Eqs. 45 and 44, the condition $n\geq r$ ensures that the last column of $\mathbf{\Theta}$ lies in the span of $\mathbf{\Theta}_{1}$ i.e.,
@@
If we pre-multiply both sides with $\mathbf{C}$, then, we can apply Eqs. 36 and 40 to reach the desired conclusion.
@@
Proposition D.3.
Suppose $\boldsymbol{\mathbf{\psi}}$ lies in the non-redundant span of $r$ distinct KEFs  and the initial condition $\boldsymbol{\mathbf{\upsilon}}_{1}$ is spectrally informative. If $(\mathbf{X},\mathbf{Y})$ is linearly consistent and the DMD problem is well-sampled, then, the columns of $\tilde{\mathbf{C}}$ are linearly independent.
@@
Proof D.4.
Firstly, if $\mathbf{A}$ is the matrix, specified by linear consistency of $(\mathbf{X},\mathbf{Y})$, that maps $\mathbf{X}$ to $\mathbf{Y}$, i.e.,
$\mathbf{A}\mathbf{X}=\mathbf{Y},$
then, we have the following identity:
We can derive this identity by using Eq. 40 to expand the data matrices $\mathbf{X}$ and $\mathbf{Y}$ as follows:
$\mathbf{A}{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\mathbf{X}}={\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\mathbf{Y}}\iff(\mathbf{A}\mathbf{C})\mathbf{\Theta}_{1}=(\mathbf{C}\mathbf{\Lambda})\mathbf{\Theta}_{1}.$
@@
$\mathbf{A}{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\mathbf{C}}\mathbf{\Theta}_{1}{}$ $=~{}\mathbf{A}\tilde{\mathbf{C}}\operatorname{diag}[\boldsymbol{\mathbf{\phi}}(\boldsymbol{\mathbf{\upsilon}}_{1})]\mathbf{\Theta}_{1},$
${\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\mathbf{C}}\mathbf{\Lambda}\mathbf{\Theta}_{1}~{}=~{}\tilde{\mathbf{C}}~{}{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\operatorname{diag}[\boldsymbol{\mathbf{\phi}}(\boldsymbol{\mathbf{\upsilon}}_{1})]\mathbf{\Lambda}}\mathbf{\Theta}_{1}{}$ $=~{}(\tilde{\mathbf{C}}\mathbf{\Lambda})~{}\operatorname{diag}[\boldsymbol{\mathbf{\phi}}(\boldsymbol{\mathbf{\upsilon}}_{1})]\mathbf{\Theta}_{1}.$
Combining these results, we have Eq. 46.
However, the matrix $\operatorname{diag}[\boldsymbol{\mathbf{\phi}}(\boldsymbol{\mathbf{\upsilon}}_{1})]\mathbf{\Theta}_{1}$ has full row rank. This can be deduced by first using $\boldsymbol{\mathbf{\upsilon}}_{1}$ being spectrally informative alongside Eq. 35 to infer that $\operatorname{diag}[\boldsymbol{\mathbf{\phi}}(\boldsymbol{\mathbf{\upsilon}}_{1})]$ is invertible. Then, we can consider the $r\times n$ Vandermonde matrix $\mathbf{\Theta}_{1}$. Since it has distinct nodes, well-sampling $(n\geq r)$ lets us use Eq. 44 to conclude that $\mathbf{\Theta}_{1}$ has linearly independent rows. Since the product of two matrices with full row rank will also have full row rank, we find that the rows of $\operatorname{diag}[\boldsymbol{\mathbf{\phi}}(\boldsymbol{\mathbf{\upsilon}}_{1})]\mathbf{\Theta}_{1}$ are linearly independent.
So, Eq. 46 becomes equivalent to the following identity:
@@
A specific interpretation of this identity, based on the construction of $\mathbf{\Lambda}$, yields the desired conclusion. To see this, observe that $\mathbf{\Lambda}$ is constructed in Eq. 39 as a diagonal matrix, by collecting the Koopman eigen-values, $\{\,\lambda_{i}\,\}_{i=1}^{r}$. Now, these eigen-values are $r$ distinct complex numbers. Hence, the resulting interpretation of Eq. 47- the columns of $\tilde{\mathbf{C}}$ are eigen-vectors of $\mathbf{A}$ corresponding to distinct eigen-values- tells us that $\tilde{\mathbf{C}}$ has linearly independent columns.
@@
D.1 Proof of Theorem 3.4
We preface the proof by establishing that $\mathbf{C}$ has full column rank. Firstly, Proposition D.3 tells us that the columns of $\tilde{\mathbf{C}}$ are linearly independent. Building on this, we need only invoke $\boldsymbol{\mathbf{\upsilon}}_{1}$ being spectrally informative alongside (35) to show that $\mathbf{C}$ also has full column rank.
DMD eigenvalues contain the Koopman eigenvalues
The key idea here is that well-sampling leads to a system of linear equations that can be peeled away, using the $r$ distinct KEFs that span our dictionary $\boldsymbol{\mathbf{\psi}}$, to produce the desired conclusion.
Firstly, well-sampling means the Companion matrix $\mathbf{T}^{*}$, defined in Eq. 7, ‘maps’ $\mathbf{X}$ to $\mathbf{Y}$ as follows:
$\mathbf{X}\mathbf{T}^{*}~{}=~{}\mathbf{Y}.$
This can be derived by considering the least-squares interpretation of Eq. 6:
@@
Since $n\geq r$, Proposition D.1 tells us that we are in the first case described above. Hence,
@@
Now, this identity can be unpacked using the Koopman mode factorization of $\mathbf{X}$ and $\mathbf{Y}$ to give the following eigenvector-eigenvalue equation: